# PDSS: A Privacy-Preserving Framework for Step-by-Step Distillation of Large Language Models

### v1

@@ 5 Conclusions
We introduced PDSS, a privacy-preserving framework for LLM distillation, addressing domain-specific knowledge privacy and resource constraints. PDSS employs a server-client architecture with prompt encoding, rationale generating, rationale decoding, and task-specific SLM training, bridging the gap between LLM and SLM while maintaining data privacy. Experiments on various text generation tasks demonstrate PDSS’s ability to enhance SLM performance with LLM support while prioritizing data privacy.
@@
Our current study faces limitations due to computational and storage constraints, which hinder our ability to experiment with larger model sizes. Additionally, our evaluation of PDSS has been restricted to the Qwen model architecture, leaving the possibility that PDSS may need to be further explored in other model architectures. We intend to tackle these issues in future research endeavors.
Appendix A Rationales Generation through COT
We utilize the rationales data generated by server-side LLM through chain-of-thought (CoT)Wei et al. (2022)Hsieh et al. (2023) technique to enhance the performance of the client’s task-specific SLM. These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models. Consider the following example: when asked “Question:A beaver is know for building prowess, their supplies come from where? Answer Choices: (a) british columbia (b) body of water (c) wooded area (d) pay debts (e) zoo”. Utilizing the chain-of-thought (CoT) technique, the LLM can generate intermediate rationales like, "The answer must be the place where beavers get their supplies. Of the above choices, only wooded areas have the supplies that beavers need.” Such rationales bridge the gap between the input and the final answer, often encapsulating valuable task-related knowledge. This knowledge would traditionally require extensive data for smaller and task-specific models to acquire. Therefore, we harness these rationales as enriched training material for small language models, employing a multi-task training paradigm that encompasses both label prediction task and rationale prediction task.
Appendix B More on Experimental Details
B.1 Hyperparameter Settings
@@
B.2 Data Splitting
For the datasets CQA/OBQA/BoolQ//ARC-E/, all splits (training, validation, and test) were downloaded from HuggingFace Lhoest et al. (2021). During the training of the Encoder-Decoder SLM, we randomly divided the training data into two equal parts. One part was designated as the public dataset, while the other part was allocated as the private dataset for the client.
B.3 Dataset Licenses
For the datasets CQA/OBQA/BoolQ//ARC-E/ were downloaded from HuggingFaceLhoest et al. (2021) and under Apache License, Version 2.0.
B.4 Machine Configuration
The experiments were conducted on machines equipped with 4 Nvidia V100 32G.
Appendix C The Definition of TokenRatio Metric
Figure 6: Comparative Analysis of Perturbed Rationales and Decoded Rationales.
Appendix D Comparative Analysis of Perturbed Rationales and Decoded Rationales
As shown in Figure 6, we conduct a comparison of the quality between the perturbed rationales and the decoded rationales, employing both the PDSS-EM and PDSS-ED methods across various privacy budgets denoted by $\epsilon$. For clarity, we designate the perturbed rationales generated using the PDSS-EM and PDSS-ED methods as P-PDSS-EM and P-PDSS-ED, respectively. Similarly, the decoded rationales derived from these methods are denoted as D-PDSS-EM and D-PDSS-ED. It’s clear that at comparable $\epsilon$ levels, the TokenRatio for decoded rationales consistently surpasses that of perturbed rationales in most tasks, when utilizing the PDSS-EM and PDSS-ED methods. This finding underscores the remarkable effectiveness of the rationales decoder module within both the PDSS-EM and PDSS-ED frameworks.

### v2

@@ 5 Conclusions
In this study, we introduce FedCoT, a federated framework designed to distill knowledge from LLMs to SLMs in resource-constrained environments. FedCoT facilitates secure knowledge transfer from LLMs to SLMs by leveraging perturbed prompts and rationales, thereby enhancing the performance of SLMs without compromising user privacy. We present two innovative privacy protection strategies, including an Adaptive Exponential Mechanism strategy, which effectively balance privacy preservation and the usability of rationales. Experiments on various text generation tasks demonstrate FedCoT’s ability to enhance SLM performance with LLM support while prioritizing data privacy.
@@
While our proposed FeCoT framework demonstrates promising results in privacy-preserving knowledge transfer from LLMs to SLMs, it is important to acknowledge several considerations that could be addressed in future work. Firstly, the framework’s performance benefits are contingent upon the server-side LLM’s CoT reasoning capabilities. Although contemporary LLMs like GPT-4 and LLaMA exhibit strong reasoning skills, frameworks such as FedCoT may encounter limitations when deployed with less sophisticated LLMs. This suggests an opportunity for further research to enhance FedCoT’s robustness against variability in LLM reasoning abilities. Secondly, our evaluation primarily focused on LLaMA and Qwen as the server-side LLMs, with client-side SLMs including Pythia, LLaMA, Qwen, and OPT. While these models are representative of current state-of-the-art architectures, extending testing to a more diverse set of LLMs could provide deeper insights into FedCoT’s generalizability.
Appendix A FedCoT’s Computational and Communication Overhead
FedCoT is designed to be efficient and scalable in resource-constrained environments. The communication overhead is minimal, with costs comparable to plaintext data transmission. Computational requirements are equivalent to standard SLM fine-tuning (SFT) on local tasks. Our experimental validation was conducted using NVIDIA V100 GPUs, demonstrating practical deployment feasibility.
Appendix B Rationales Generation through CoT
We utilize the rationales data generated by server-side LLM through chain-of-thought (CoT)wei2022chainhsieh2023distilling technique to enhance the performance of the client’s task-specific SLM. These rationales justify the predicted labels and serve as insightful guidance for training smaller and domain-specific models. Consider the following example: when asked “Question:A beaver is know for building prowess, their supplies come from where? Answer Choices: (a) british columbia (b) body of water (c) wooded area (d) pay debts (e) zoo”. Utilizing the chain-of-thought (CoT) technique, the LLM can generate intermediate rationales like, "The answer must be the place where beavers get their supplies. Of the above choices, only wooded areas have the supplies that beavers need.” Such rationales bridge the gap between the input and the final answer, often encapsulating valuable task-related knowledge. This knowledge would traditionally require extensive data for smaller and task-specific models to acquire. Therefore, we harness these rationales as enriched training material for small language models, employing a multi-task training paradigm that encompasses both label prediction task and rationale prediction task.
Appendix C More on Experimental Details
C.1 Hyperparameter Settings
@@
C.2 Data Splitting
For the datasets CQA/OBQA/BoolQ//ArcE/, all splits (training, validation, and test) were downloaded from HuggingFace lhoest-etal-2021-datasets. During the training of the Encoder-Decoder SLM, we randomly divided the training data into two equal parts. One part was designated as the public dataset, while the other part was allocated as the private dataset for the client.
C.3 Dataset Licenses
For the datasets CQA/OBQA/BoolQ//ARC-E/ were downloaded from HuggingFacelhoest-etal-2021-datasets and under Apache License, Version 2.0.
C.4 Machine Configuration
The experiments were conducted on machines equipped with 4 and 8 NVIDIA V100 32G.